OpenStudy (anonymous):

What number would you have to add to both sides to complete the square? x2 + 12x = –6 A. –36 B. 36 C. 144 D. –144 I thinks it's c but I'm not sure

OpenStudy (johnweldon1993):

So to complete the square ...you take the coefficient of the 'x' ...divide it by 2....and then square the result... so that coefficient is 12... half of that is 6 6 square = 36 ....so we would add 36 to both sides of our equation

OpenStudy (anonymous):

omg thank you soo much

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (anonymous):

can you help with more?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (anonymous):

Solve by using the perfect squares method. x2 + 16x + 64= 0 A. –8 B. –16 C. 16 D. 8 i think b

OpenStudy (anonymous):

sorry i meant a

OpenStudy (johnweldon1993):

Alright so perfect squares method means... you can turn something like that into something like (x + 8)^2 Why? because when you factor that out you get (x + 8)(x + 8) x^2 + 8x + 8x + 64 x^2 + 16x + 64 Well that is correct...so it looks like my hunch was correct....now what would make (x + 8)^2 = 0? Looks like if you add -8 to 8..you get 0 right? So -8 it is....you are correct :)

OpenStudy (anonymous):

yay! Solve. x2 – 5x + 6 = 0 A. {3, –2} B. {2, 3} C. {–2, –3} D. {5, –1} is it b

OpenStudy (johnweldon1993):

Indeed it is B See you're a pro at these :)

OpenStudy (anonymous):

Solve. 2x2 + x – 1 = 9 i think it is 2, -5/2

OpenStudy (johnweldon1993):

You would be correct!

OpenStudy (anonymous):

yay! 2x2 + 8x = 0 A. {0, 4} B. (0,-1/4 C. {0, –4} D. {3, 5}

OpenStudy (johnweldon1993):

And your guess?

OpenStudy (anonymous):

c

OpenStudy (johnweldon1993):

Correct :)

OpenStudy (anonymous):

okay last one i swear Use the discriminant to determine the number and type of solutions the equation has. x2 + 6x + 12 = 0 A. two rational solutions B. two irrational solutions C. no real solution D. one real solution is it b?

OpenStudy (anonymous):

Thank you soo much I really appreciate you helping me @johnweldon1993

OpenStudy (johnweldon1993):

Not quite...irrational just means a decimal number that will go on forever But the discriminant is \[\large b^2 - 4ac\] \[\large 6^2 - 4(1)(12)\] \[\large 36- 48 = -12\] When the discriminate comes out as a negative...it shows that you have no REAL solutions to the equation...you instead have a complex solution...so this would be C

OpenStudy (johnweldon1993):

And yes of course @sierraleone17 :)

OpenStudy (anonymous):

thank you thank you thank you

OpenStudy (johnweldon1993):

Anytime :)